The paragraph starting at page 18, line -5 of the published version, or at page 13, line -9 of this version (beginning “This establishes that the set of adelic points”) should be replaced with the following:

This establishes that all adelic points at which the sum of the local invariants of both E_{1} and E_{2} +C is 0 is contained in the set of points whose local invariant at 1033 is (1/2, 0, 1/2). The new phenomenon is that at 1033, adelic points lying over the singular 𝔽_{1033}-point [271 : 795 : 1] have local invariants (1/2, 0, 1/2), while other points go to (0, 0, 0). For the second statement, it is sufficient to check one point on S over ℚ_{1033} whose reduction mod 1033 is smooth; for the first, we consider all lifts mod 1033^{2} and verify that the only one for which the local invariant cannot be concluded to be (1/2, 0, 1/2) does not lift mod 1033^{3}. A short search over points of ℙ^{2} reducing to [271 : 795 : 1] mod 1033 then finds the smallest rational point [55782, −1298, 126, −2978].